Solve by substitution method
5x + 7y =10
x= 5-2y
What is the solution pair

Answer by Alan3354(69443)

Solve by substitution method
5x + 7y =10
x= 5-2y
----------
Sub for x in 1st eqn
5(5-2y) + 7y = 10
25-10y + 7y = 10
-3y = -15
y = 5
----------
x = -5
--> (-5,5)